# A Comprehensive Survey with Critical Analysis for Deepfake Speech Detection

### v1

@@ I INTRODUCTION
Focusing on human speech, this paper provides a comprehensive survey for the task of Deepfake Speech Detection (DSD). To this end, the milestones presenting the development progress of the DSD task are first presented in Fig. 1. As the figure showns, the earliest public dataset and challenge proposed for the DSD task was introduced in 2015, focusing exclusively on the English language. Then, the first challenge for video deepfake detection (DFDC [15]) was introduced in 2020. In subsequent years, datasets for the DSD task in Japanese [16], Korean [16], and Chinese [17] were introduced in 2021 and 2022, respectively. Recently, in 2024, multilingual datasets for the DSD task have been published, including MLAAD [18] for conversational speech and SVDD [19] for singing. Fig. 1 also highlights a growing number of papers, datasets, and challenge competitions for the DSD task from 2021 to the present. This trend indicates that the DSD task has recently gained prominence and has attracted significant interest from the research community.
@@
To further understand the DSD task, we summarized recent survey papers related to the DSD task in Table I. As shown in the table, most of these surveys focus on detecting general fake data (e.g., images, videos, audio, or text), with audio or human speech typically being addressed only as a subsection or a part of the broader discussion [8, 2, 3]. Therefore, the main techniques, existing concerns, and potential research for the DSD task have not been comprehensively analyzed in these papers. Among the survey papers, only two survey papers of [5] and [9] focus on the DSD task. However, as conventional surveys, these papers primarily summarize the technologies used to construct a DSD system such as datasets, feature extraction, classification model, loss functions, rather than providing a comprehensive analysis and highlighting existing concerns. For instance, while challenge competitions proposed for the DSD task are very important in advancing the research community, their importance and various aspects have not been thoroughly analyzed (i.e., The number of research teams participating in these competitions and their results are interesting to analyzed, etc). Although this information reflects the level of interest in DSD within the research community, it has not been addressed in any existing survey papers. The second concern is related to public datasets proposed for the DSD task. In particular, the current survey papers do not adequately analyze the imbalance among (1) the number of utterances, (2) the AI-synthesized speech systems used to generate fake speech; and (3) the original/real human speech resource used to generate fake speech utterances. These key factors are essential in creating a high-quality DSD dataset for evaluating DSD models. Additionally, survey papers are at risk of becoming outdated as new datasets, techniques, and models continue to emerge. However, current surveys do not offer solutions for regularly updating essential information, such as details about challenge competitions, public datasets, and the top-performing models on specific datasets.
@@
As regards technologies used to construct a DSD model such as feature extraction, classification model, or loss functions, current survey papers mainly summarize and then present conclusions rather than conducting experiments to provide strong evidence and validation.
Given the existing concerns about the current survey papers for the DSD task, these motivate and inspire us to provide a much more comprehensive survey in this paper. By addressing these concerns, we mainly contribute:
  • We provide a comprehensive analysis and then indicate concerns related to three main topics: The current challenge competition, the published datasets, and the deep-learning-based techniques used to develop a DSD system. We construct each topic by describing three main parts including ‘Analysis’, ‘Discussion’, and ‘Contribution’. The ‘Analysis’ summarizes concrete information about the topic. Meanwhile, the ‘Discussion’ indicates concerns in each topic. Finally, the ‘Contribution’ provides our suggestion and solution to further contribute to each topic.
  • To solve the out-of-date issue of a survey paper, we set up a Github repository to update further challenge competitions, public datasets, and top-performance systems.
  • More than a survey, we conduct extensive experiments to verify assumptions from the comprehensive analysis, achieving a competitive DSD model. Given the analysis and experimental results, we indicate potential research directions for the DSD task.
@@ II Challenge Competitions Proposed for Deepfake Speech Detection
Analysis: Challenge competitions for the DSD task play a crucial role in motivating the research community. These competitions not only introduce new benchmark datasets but also host workshops where research teams can discuss their ideas and share their motivations. This environment encourages the community to publish more datasets and develop new techniques to address the DSD challenges. To analyze DSD challenge competitions, we first summarize all challenges in Table II. Importantly, we will continuously update information about future DSD challenge competitions in our GitHub repository [^1] https://github.com/AI-ResearchGroup/A-Comprehensive-Survey-with-Critical-Analysis-for-Deepfake-Speech-Detection.
@@
DSD challenge competitions, which explore fake speech in a conversation, can be separated into two groups. The first group is proposed for only audio [20, 29, 21, 23, 27, 22, 17, 24]. Meanwhile, the second group is for video in which a fake video is identified by fake audio, fake image, or both fake audio and image [15, 26]. This indicates that DSD is not only treated as an individual task independently but also considered as a sub-task in multimodal systems. It is also evident that the second group, which focused on fake video detection, has attracted significantly more research teams (e.g. 2,114 teams in the DFDC challenge [15]) compared to the first group (e.g. the largest team count was 54 in the ASVspoof 2021 challenge [23]). This provides an insight that fake video detection is a more compelling task, drawing greater interest and participation from research teams. Regarding top-1 systems in these challenge competitions, they leveraged the ensemble techniques which combine a wide range of input features or multiple models (i.e., most submitted systems mainly use deep learning based models).
@@
Discussion: Given the recent analysis of challenge competitions proposed for the DSD task, some concerns can be indicated. Firstly, the DSD task has drawn attention from the research community and is now recognized as one of the critical components in a complex system of deepfake detection. However, most current challenge competitions are limited to single languages, such as Chinese or English, and primarily focus on detecting fake speech within conversations. Second, some challenge competitions have not published datasets for different reasons. For example, FTC [22] was organized by the US government, and the top-performing systems are used by the US government. Similarly, ADD 2023 [24] only provides the dataset for the teams that attended during the competition. These limitations hinder research motivation and further development once the challenges conclude. Third, it is recognized that fake speech utterances are mainly generated from deep-learning-based speech generation systems. Therefore, if selected deep-learning-based speech generators are not general or up-to-date, this significantly affects the effectiveness and quality of the challenge competition. This highlights the need for collaboration between two tasks of deep-learning-based speech generation and detection within the same challenge competition. Competitions like ASVspoof 2024 [27] and ADD 2022 [17] have addressed this by not only published datasets but also presented a two-phase or two-track challenge in which the first phase/track is for Deepfake Speech Generation and the second phase/track is for Deepfake Speech Detection. Finally, regarding techniques used in these competitions, ensemble models have become widely used to enhance performance in many challenge competitions, often leading research teams to develop top-performing systems. However, this approach has several drawbacks, including limited explainability, increased system complexity, high training costs, and concerns related to power consumption and green AI. Therefore, different aspects of using deep-learning-based models such as using a single model, low complexity, or real-time inference can be regarded as main constraints in challenge competitions for the DSD task in the future. For example, the DCASE challenge Task 1 [49] for Sound Scene Classification requires the submitted systems to obey two constraints: (1) not larger than 128 K parameters and (2) not larger than 30 MMAC units.
@@
  • We present and highlight the important role of DSD challenge competitions. We then provide a comprehensive analysis and indicate the concerns.
  • We continue updating new challenge competitions in the future by creating a Github project [^2] https://github.com/AI-ResearchGroup/A-Comprehensive-Survey-with-Critical-Analysis-for-Deepfake-Speech-Detection. The GitHub repository serves as a reference for up-to-date information on challenge competitions and current concerns. In other words, it provides a summary of challenge competitions related to the DSD task, ensuring that this survey paper remains current.
@@ III Public Datasets Proposed For Deepfake Speech Detection
As illustrated in Fig.3, the number of public datasets for the DSD task has grown significantly in recent years. Most of these datasets include both clean and noisy speech. Notably, nearly all datasets have been designed for English, with WaveFake [16], KoDF [50], and ADD 2022 [17] being the exceptions, focusing on Japanese, Korean, and Chinese languages, respectively. Recently, the first multilingual datasets for the DSD task were introduced in [47] and [19]. The MLAAD dataset [18] provides fake speech in conversations generated in 23 widely spoken languages. Meanwhile, the SVDDD dataset [19] was proposed for deepfake singing detection with six different languages (i.e., the Chinese songs are the majority).
@@
Most deepfake datasets are generated from one of three generator techniques: Text-to-Speech (TTS), Voice Conversion (VC), and Adversarial Attacks (AT), as shown in Table IV. Notably, ASVspoof 2024 [27] is the first dataset that uses AT systems to generate fake speech. While TTS systems generate fake speech from text, VC systems generate fake speech from real speech (e.g. audio). To mimic the target speakers, TTS and VC systems attempt to explore the audio embeddings extracted from the target speakers. The audio embeddings extracted from target speakers are treated as a part of the feature map in the entire network architecture in TTS and VC systems. Regarding AC systems, they mainly apply Malafide [124] and Malocopula [125] methods to generate fake speech. Both Malafide [124] and Malocopula [125] methods involve leveraging filter banks. Malafide [124] applies multiple techniques of linear time-invariant (LTI), non-causal filter, and the coefficients (e.g., tap weights) to create TTS/VC-based fake speech that mimics the target speaker. Meanwhile, Malocopula [125] combines both linear filter and non-linear filter (e.g., one-dimensional convolutional layer) to replicate the target speaker’s voice.
@@
To compare among DSD datasets, we analyze three different aspects: (1) the number of fake utterances; (2) the AI-synthesized speech systems used to generate fake speech; and (3) the original/real human speech resource used to generate fake speech utterances. As Table III shows, most datasets present lower than 300,000 utterances of fake speech, expected ASVspoof 2021 (DF Task) [23], ASVspoof 2024 [27], AV-Deepfake1M dataset [25, 26], and DFDC dataset [15] with 589212, 815262, 860039, and 104500 fake samples, respectively. Although DFDC [15, 82] and AV-Deepfake1M dataset [25, 26] present a large number of fake data, this was proposed for video in which audio may not be fake. Additionally, these fake utterances were generated from only a few deep-learning-based speech-generation systems. Indeed, two TTS models of VITS [97], YoursTTS [98] and one TTS model [82] were used to generate fake speech in DFDC [15] and AV-Deepfake1M dataset [25, 26] datasets, respectively. On the other hand, the ASVspoof 2021 (DF Eva) dataset [23] contains 589212 fake utterances, generated using over 100 voice conversion (VC) and text-to-speech (TTS) systems. To catch up with state-of-the-art deepfake speech generators, Table IV presents the architectures and resources of deepfake speech generators. The table indicates that the ASVspoofing series shows up-to-date and diverse deepfake speech generators compared to the others. In terms of the original human speech resources, most DSD datasets are based on recordings from a limited number of speaker volunteers. For example, although the ASVspoof 2021 (DF Eva) dataset [23] used 100 VC and TTS systems to create fake utterances, the real speech resource is from 107 speaker volunteers. Some DSD datasets of AV-Deepfake1M [25, 26], CFAD [43] leveraged the large and available human speech datasets to generate fake utterances such as Voxceleb2 [37], AISHELLI-3 [35], MAGICDATA [46]. However, these datasets use a limited number of speech generators (e.g. 2 TTS and 11 TTS for AV-Deepfake1M [25, 26] and CFAD [43], respectively).
@@
Secondly, we also highlight an imbalance among DSD datasets regarding three aspects: (1) the number of fake utterances; (2) the AI-synthesized speech systems used to generate fake speech; and (3) the original/real human speech resource used to generate fake speech utterances. The imbalance can be clearly described in Fig. 5 regarding DSD datasets using speaker volunteers.
  • The number of utterances: The quantity of utterances within the datasets is not uniform. Some datasets may contain a large number of samples, while others have significantly fewer. A small number of real or fake utterances within datasets (e.g. FakeAVCeleb [36], ADD [17]) limits the model’s exposure to a wide variety of speech patterns and scenarios, affecting the detection robustness and generalization new, unseen data. Additionally, a controlled ratio between real and fake samples created within datasets (e.g ASVspoof 2024 [27], ASVsproof 2021 [23]) also ensure diversity of fake techniques and avoid overfitting on the fake data, especially if the fake samples are generated using similar techniques. Therefore, maintaining a moderately controlled ratio between real and fake utterances, along with a diverse range of these utterances, is essential for future dataset development.
@@
  • Deepfake speech generation systems: The variety of deep-learning-based systems used to generate deepfake speech is another area of concern. As Table IV shows, some of datasets such as MLAAD [47], ASVspoof 2021(DF task) [23], ASVspoof 2024 [27] present more than 20 systems (e.g. TTS, VC, or AT systems). Among these datasets, ASVspoof 2021 (DF Task) [23] and ASVspoof 2024 [27] present diverse TTS, VC, and AT systems. In particular, while more than 100 TTS and VC are for ASVspoof 2021 (DF Task) [23], 28 TTS, VC, and AT are used in ASVspoof 2024 [27]. Although MLAAD [47] has been the unique multiple-language dataset currently, fake speech in this dataset was only generated from TTS systems (e.g. 54 TTS systems). Overall, some datasets may predominantly feature speech synthesized by a few specific deep-learning-based generators or techniques, while others might include a broader range. Datasets generated from a limited number of deep-learning-based generators possibly lead to over-specialization, reducing the model’s ability to detect deepfakes generated by other systems and affecting the performance in real-world scenarios. Therefore, this imbalance motivates the research community to create more diverse datasets that include a wide range of AI-synthesized speech methods.
@@
  • Real human speech resource: The source of real voice plays a crucial role in shaping the effectiveness, generalizability, and ethical aspects of deepfake detection models. As highlighted in Table IV, there are two main sources for building DSD datasets: voice samples from volunteer speakers or from existing datasets. Voice samples from volunteers offer greater control over diversity (if managed thoroughly) and address ethical concerns, as they are collected with explicit informed consent. However, this approach can be resource-intensive in terms of time and cost and may not scale efficiently. In contrast, utilizing existing human speech datasets offers better accessibility and scalability. However, it may introduce biases toward certain groups, such as public figures, reducing diversity in real-world applications and especially raising significant ethical issues. These problems suggest other balanced approaches to build DSD datasets that consider both diversity and scalability in the future.
@@
Our contribution: Given the analysis and the discussion above, our work mainly contributes:

### v4

@@ I INTRODUCTION
Focusing on human speech, this paper provides a comprehensive survey for the task of Deepfake Speech Detection (DSD). To this end, the milestones presenting the development progress of the DSD task are first presented in Fig. 1. As the figure shows, the earliest public dataset and challenge proposed for the DSD task was introduced in 2015, focusing exclusively on the English language. Then, the first challenge for video deepfake detection (DFDC [15]) was introduced in 2020. In subsequent years, datasets for the DSD task in Japanese [16], Korean [16], and Chinese [17] were introduced in 2021 and 2022, respectively. Recently, in 2024, multilingual datasets for the DSD task have been published, including MLAAD [18] for conversational speech and SVDD [19] for singing. Fig. 1 also highlights a growing number of papers, datasets, and challenge competitions for the DSD task from 2021 to the present. This trend indicates that the DSD task has recently gained prominence and has attracted significant interest from the research community.
@@
To further understand the DSD task, we summarized recent survey papers related to the DSD task in Table I. As shown in the table, most of these surveys focus on detecting general fake data (e.g., images, videos, audio, or text), with audio or human speech typically being addressed only as a subsection or a part of the broader discussion [8, 2, 3]. Therefore, the main techniques, existing concerns, and potential research for the DSD task have not been comprehensively analyzed in these papers. Among the survey papers, only two survey papers of [5] and [9] focus on the DSD task. However, as conventional surveys, these papers primarily summarize the technologies used to construct a DSD system such as datasets, feature extraction, classification model, loss functions, rather than providing a comprehensive analysis and highlighting existing concerns. For instance, while challenge competitions proposed for the DSD task are very important in advancing the research community, their importance and various aspects have not been thoroughly analyzed (e.g., the number of research teams participating in these competitions and their results are interesting to analyze). Although this information reflects the level of interest in DSD within the research community, it has not been addressed in any existing survey papers. The second concern is related to public datasets proposed for the DSD task. In particular, the current survey papers do not adequately analyze the imbalance among (1) the number of utterances, (2) the AI-synthesized speech systems used to generate fake speech; and (3) the original/real human speech resource used to generate fake speech utterances. These key factors are essential in creating a high-quality DSD dataset for evaluating DSD models. Additionally, survey papers are at risk of becoming outdated as new datasets, techniques, and models continue to emerge. However, current surveys do not offer solutions for regularly updating essential information, such as details about challenge competitions, public datasets, and the top-performing models on specific datasets.
@@
Regarding technologies used to construct a DSD model such as feature extraction, classification model, or loss functions, current survey papers mainly summarize and then present conclusions rather than conducting experiments to provide strong evidence and validation.
The above concerns about the existing survey papers for the DSD task motivate and inspire us to provide a much more comprehensive survey in this paper. By addressing these concerns, we make the following main contributions:
  • We provide a comprehensive analysis and then indicate concerns related to three main topics: The current challenge competition, the published datasets, and the deep-learning-based techniques used to develop a DSD system. Each topic consists of three main parts: ‘Analysis’, ‘Discussion’, and ‘Contribution’. The ‘Analysis’ summarizes concrete information about the topic. The ‘Discussion’ indicates concerns in each topic. Finally, the ‘Contribution’ provides our suggestion and solution to further improve each topic.
  • To solve the out-of-date issue of a survey paper, we set up a Github repository to update further challenge competitions, public datasets, and top-performance systems. New versions of the paper are also continually updated on ‘https://arxiv.org’.
  • More than a survey, we conduct extensive experiments to verify assumptions from the comprehensive analysis (i.e., different types of data augmentation, multiple input features, multiple network architectures, cross-dataset and cross-language evaluation, etc.), achieving a competitive DSD model. Given the analysis and experimental results, we indicate potential research directions for the DSD task.
@@ II Challenge Competitions Proposed for Deepfake Speech Detection
Analysis: Challenge competitions for the DSD task play a crucial role in motivating the research community. These competitions not only introduce new benchmark datasets but also host workshops where research teams can discuss their ideas and share their motivations. This environment encourages the community to publish more datasets and develop new techniques to address the DSD challenges. To analyze DSD challenge competitions, we first summarize all challenges in Table II. Importantly, we will continually update information about future DSD challenge competitions in our GitHub repository [^1] https://github.com/AI-ResearchGroup/A-Comprehensive-Survey-with-Critical-Analysis-for-Deepfake-Speech-Detection.
@@
DSD challenge competitions, which explore fake speech in a conversation, can be separated into two groups. The first group is proposed for only audio [20, 29, 21, 23, 27, 22, 17, 24]. Meanwhile, the second group is for video in which a fake video is identified by fake audio, fake image, or both fake audio and image [15, 26]. This indicates that DSD is not only treated as an individual task independently but also considered as a sub-task in multimodal systems. It is also evident that the second group, which focuses on fake video detection, has attracted significantly more research teams (e.g., 2,114 teams in the DFDC challenge [15]) compared to the first group (e.g., the largest team count was 74 in the ASVspoof 2021 challenge [23]). This provides an insight that fake video detection is a more compelling task, drawing greater interest and participation from research teams. Regarding top-1 systems in these challenge competitions, they leveraged the ensemble techniques which combine a wide range of input features or multiple models (i.e., most submitted systems mainly use deep learning based models).
@@
Discussion: Given the recent analysis of challenge competitions proposed for the DSD task, some concerns can be indicated. Firstly, the DSD task has drawn attention from the research community and is now recognized as one of the critical components in a complex system of deepfake detection. However, most current challenge competitions are limited to single languages, such as Chinese or English, and primarily focus on detecting fake speech within conversations. Secondly, some challenge competitions have not published datasets for different reasons. For example, FTC [22] was organized by the US government, and the top-performing systems are used by the US government. Similarly, ADD 2023 [24] only provides the dataset for the teams that attended during the competition. These limitations hinder research motivation and further development once the challenges conclude. Third, it is recognized that fake speech utterances are mainly generated from deep-learning-based speech generation systems. Therefore, if selected deep-learning-based speech generators are not general or up-to-date, this significantly affects the effectiveness and quality of the challenge competition. This highlights the need for collaboration between two tasks of deep-learning-based speech generation and detection within the same challenge competition. Competitions like ASVspoof 2024 [27] and ADD 2022 [17] have addressed this by not only published datasets but also presented a two-phase or two-track challenge in which the first phase/track is for Deepfake Speech Generation and the second one is for Deepfake Speech Detection. Finally, regarding techniques used in these competitions, ensemble models have become widely leveraged to enhance performance in many challenge competitions, enabling research teams to develop top-performing systems. However, this approach has several drawbacks, including limited interpretability, increased system complexity, high training costs, and concerns related to power consumption and green AI. Therefore, different aspects of using deep-learning-based models such as using a single model, low complexity, or real-time inference can be regarded as main constraints in challenge competitions for the DSD task in the future. For example, the DCASE challenge Task 1 [49] for Sound Scene Classification requires the submitted systems to obey two constraints: (1) not larger than 128 K parameters and (2) not larger than 30 MMAC units.
@@
  • We present and highlight the important role of DSD challenge competitions. We then provide a comprehensive analysis and indicate the existing concerns.
  • We continue updating new challenge competitions in the future by creating a Github project [^2] https://github.com/AI-ResearchGroup/A-Comprehensive-Survey-with-Critical-Analysis-for-Deepfake-Speech-Detection. The GitHub repository serves as a reference for up-to-date information on challenge competitions and current concerns. In other words, it provides a summary of challenge competitions related to the DSD task, ensuring that this survey paper stays updated.
@@ III Public Datasets Proposed For Deepfake Speech Detection
As illustrated in Fig.3, the number of public datasets for the DSD task has grown significantly in recent years. Most of these datasets include both clean and noisy speech. Notably, nearly all datasets have been designed for English, with WaveFake [16], KoDF [50], and ADD 2022 [17] being the exceptions, focusing on Japanese, Korean, and Chinese languages, respectively. Recently, the first multilingual datasets for the DSD task were introduced in [47] and [19]. The MLAAD dataset [18] provides fake speech in conversations generated in 23 widely spoken languages. Meanwhile, the SVDD dataset [19] was proposed for deepfake singing detection with six different languages (i.e., the Chinese songs are the majority).
@@
Most deepfake datasets are generated from one of three generator techniques: Text-to-Speech (TTS), Voice Conversion (VC), and Adversarial Attacks (AT), as shown in Table IV. Notably, ASVspoof 2024 [27] is the first dataset that uses AT systems to generate fake speech. While TTS systems generate fake speech from text, VC systems generate fake speech from real speech (e.g., audio). To mimic the target speakers, TTS and VC systems attempt to explore the audio embeddings extracted from the target speakers. These audio embeddings are treated as a part of the feature map in the entire network architecture in TTS and VC systems. Regarding AT systems, they mainly apply Malafide [124] and Malocopula [125] methods to generate fake speech. Both Malafide [124] and Malocopula [125] methods involve leveraging filter banks. Malafide [124] applies multiple techniques of linear time-invariant (LTI), non-causal filter, and the coefficients (e.g., tap weights) to create TTS/VC-based fake speech that mimics the target speaker. Meanwhile, Malocopula [125] combines both linear filter and non-linear filter (e.g., one-dimensional convolutional layer) to replicate the target speaker’s voice.
@@
To compare among DSD datasets, we analyze three different aspects: (1) the number of fake utterances; (2) the AI-synthesized speech systems used to generate fake speech; and (3) the original/real human speech resource used to generate fake speech utterances. As Table III shows, most datasets present lower than 300,000 utterances of fake speech, except ASVspoof 2021 (DF Task) [23], ASVspoof 2024 [27], and AV-Deepfake1M dataset [25, 26] with 589212, 815262, and 860039 fake samples, respectively. Although DFDC [15, 82] and AV-Deepfake1M dataset [25, 26] present a large number of fake data, this was proposed for video in which audio may not be fake. Additionally, these fake utterances were generated from only a few deep-learning-based speech-generation systems. Indeed, two TTS models of VITS [97], YoursTTS [98] and one TTS model [82] were used to generate fake speech in DFDC [15] and AV-Deepfake1M dataset [25, 26] datasets, respectively. On the other hand, the ASVspoof 2021 (DF Eva) dataset [23] contains 589212 fake utterances, generated using over 100 voice conversion (VC) and text-to-speech (TTS) systems. To catch up with state-of-the-art deepfake speech generators, Table IV presents the architectures and resources of deepfake speech generators. The table indicates that the ASVspoofing series show up-to-date and diverse deepfake speech generators compared to the others. In terms of the original human speech resources, most DSD datasets are based on recordings from a limited number of speaker volunteers. For example, although the ASVspoof 2021 (DF Eva) dataset [23] used 100 VC and TTS systems to create fake utterances, the real speech resource is from 107 speaker volunteers. Some DSD datasets of AV-Deepfake1M [25, 26], CFAD [43] leveraged the large and available human speech datasets to generate fake utterances such as Voxceleb2 [37], AISHELLI-3 [35], MAGICDATA [46]. However, these datasets use a limited number of speech generators (e.g., 2 TTS and 11 TTS for AV-Deepfake1M [25, 26] and CFAD [43], respectively).
@@
Secondly, another limitation of currently available datasets is that they focus on a limited number of DSD use cases. In particular, two use cases should be clearly distinguished: 1) detecting deepfakes without access to the original voice, and 2) detecting deepfakes with access to the original voice. The current datasets are designed for addressing the former but not the latter use case as they lack authentic-cloned speech pairing. Another highly relevant use case that should be addressed in the future is partially deepfake speech whereby just a part of the speech is being replaced by a synthetic component. Thirdly, we highlight an imbalance among DSD datasets regarding three aspects: (1) the number of fake utterances; (2) the AI-synthesized speech systems used to generate fake speech; and (3) the original/real human speech resource used to generate fake speech utterances. The imbalance can be clearly described in Fig. 5 regarding DSD datasets using speaker volunteers.
  • The number of utterances: The quantity of utterances within the datasets is not uniform. Some datasets may contain a large number of samples, while others have significantly fewer. A small number of real or fake utterances within datasets (e.g., FakeAVCeleb [36], ADD [17]) limits the model’s exposure to a wide variety of speech patterns and scenarios, affecting the detection robustness and generalization on new, unseen data. Additionally, a controlled ratio between real and fake samples created within datasets (e.g., ASVspoof 2024 [27], ASVsproof 2021 [23]) also ensure diversity of fake techniques and avoid overfitting on the fake data, especially if the fake samples are generated using similar techniques. Therefore, maintaining a moderately controlled ratio between real and fake utterances, along with a diverse range of these utterances, is essential for future dataset development.
@@
  • Deepfake speech generation systems: The variety of deep-learning-based systems used to generate deepfake speech is another area of concern. As Table IV shows, some of datasets such as MLAAD [47], ASVspoof 2021(DF task) [23], ASVspoof 2024 [27] present more than 20 systems (e.g., TTS, VC, or AT systems). Among these datasets, ASVspoof 2021 (DF Task) [23] and ASVspoof 2024 [27] present diverse TTS, VC, and AT systems. In particular, while more than 100 TTS and VC are for ASVspoof 2021 (DF Task) [23], 28 TTS, VC, and AT are used in ASVspoof 2024 [27]. Although MLAAD [47] has been the unique multiple-language dataset currently, fake speech in this dataset was only generated from 54 TTS systems. Overall, some datasets may predominantly feature speech synthesized by a few specific deep-learning-based generators or techniques, while others might include a broader range. Datasets generated from a limited number of deep-learning-based generators possibly lead to over-specialization, reducing the model’s ability to detect deepfakes generated by other systems and affecting the performance in real-world scenarios. Therefore, this imbalance motivates the research community to create more diverse datasets that include a wide range of AI-synthesized speech methods.
@@
  • Real human speech resource: The source of real voice plays a crucial role in shaping the effectiveness, generalization, and ethical aspects of deepfake detection models. As highlighted in Table IV, there are two main sources for building DSD datasets: voice samples from volunteer speakers or from existing datasets. Voice samples from volunteers offer greater control over diversity (if managed thoroughly) and address ethical concerns, as they are collected with explicitly informed consent. However, this approach can be resource-intensive in terms of time and cost and may not scale efficiently. In contrast, utilizing existing human speech datasets offers better accessibility and scalability. However, it may introduce biases toward certain groups, such as public figures, reducing diversity in real-world applications and especially raising significant ethical issues. These problems suggest other balanced approaches to build DSD datasets that consider both diversity and scalability in the future.
@@
Our contribution: Given the analysis and the discussion above, we make the following main contributions: